the amount (A) on deposit in a bank account bearing simple interest is given by the formula
A=P+PRT
find A when P=$1200
R=0.08
T=3

Answer by robertb(5830)

A = P(1+RT) = 1200(1 + 0.24) = 1200*1.24 = 1488 $